# TriNeRFLet: A Wavelet Based Multiscale Triplane NeRF Representation

### v1

@@ Abstract
In recent years, the neural radiance field (NeRF) model has gained popularity due to its ability to recover complex 3D scenes. Following its success, many approaches proposed different NeRF representations in order to further improve both runtime and performance. One such example is Triplane, in which NeRF is represented using three 2D feature planes. This enables easily using existing 2D neural networks in this framework, e.g., to generate the three planes. Despite its advantage, the triplane representation lagged behind in its 3D recovery quality compared to NeRF solutions. In this work, we propose TriNeRFLet, a 2D wavelet-based multiscale triplane representation for NeRF, which closes the 3D recovery performance gap and is competitive with current state-of-the-art methods. Building upon the triplane framework, we also propose a novel super-resolution (SR) technique that combines a diffusion model with TriNeRFLet for improving NeRF resolution.
@@ 1 Introduction
Figure 1: Our approach improves the quality of NeRF reconstruction. (Left) Using a multiscale wavelet representation with Triplane improves 3D reconstruction quality; (Right) Combining diffusion-based super-resolution with TriNeRFLet leads to high-resolution reconstruction from low-resolution 2D views.
@@
While being useful due to its special 2D structure, the reconstruction quality achieved by Triplane lagged behind other efficient multiview reconstruction methods such as instang NGP (INGP) [28], which is an improvement of NeRF, and 3D Gaussian splatting [20]. Due to its structure, only Triplane entries that are included in the training views rays are learned. Thus, there might be entries at the planes that will still have their initial random values also after training finishes. These random features will be used by novel views that use them and therefore will deteriorate the quality of their generation and create artifacts in them.
@@
To tackle these drawbacks, we present a novel Triplane representation that relies on a multiscale 2D wavelet structure. Instead of learning the feature planes directly, we learn the wavelet features of several resolutions, while regularizing the wavelet representation to be sparse. Thus, regions that are covered by the training views are learned in a similar way to the case of a regular Triplane, and regions that are not covered by the training views are updated by a lower resolution estimate according to how much they are affected by their neighboring regions. Figure 1(left) shows the improvement in reconstruction quality attained by this representation. Figure 2 illustrates how regions on the plane may affect each other due to the wavelet representation.
@@
We use $L_{1}$ regularization on the wavelet coefficients. The goal is to sparsify the coefficients that are not trained by the given views such that only the coefficients that were trained will be used. Our proposed structure is inspired by the wavelet literature, which indicates that wavelet features can represent data accurately while being sparse [11].
@@
To further improve the training, we also perform the training in a multi-scale fashion, starting with a lower resolution version of the images that updates only the coarse coefficients of the wavelet. Then, after some iterations we increase back the resolution and add more resolution to the wavelet representation.
@@
Our proposed TriNeRFLet solution closes the performance gap of Triplane against other multiview 3D reconstruction methods and makes it competitive with current state-of-the-art (SOTA) methods. In addition, we show how we may use pre-trained diffusion models for 2D super-resolution (SR) [33, 34] with our TriNeRFLet representation in order to propose a novel NeRF SR approach. Figure 1(Right) demonstrates the improvement this approach attains. We show on various experiments the advantage of our TriNeRFLet in novel view reconstruction compared to INGP and 3D Gaussian splatting and the superiority of our SR solution compared to competing approaches that do not require dedicated training for multiview or 3D data.
@@
Figure 2: The multiscale wavelet representation transfers information between points on the triplane planes: Although the Red dot on the plane was not covered by the training images views, it will get a coarser estimate from a lower resolution wavelet layer that corresponds also to the blue dot, which was updated in training.
@@
(b) TriNeRFLet training scheme. First, wavelet features are transformed into the Triplane domain. Next, pixels are rendered using these features in order to fit them to their ground-truth values. The high frequencies channels of LH, HL and HH from all wavelet levels get regularized by $L_{1}$ loss.
(c) In the wavelet domain, the low-frequency coefficients $\textrm{LL}_{N_{base}}$ are observed at the top left corner (the bright spot). The rest are the high frequencies channels $\textrm{LH}_{N_{base}}$, $\textrm{HL}_{N_{base}}$ and $\textrm{HH}_{N_{base}}$. All the other LH, HL and HH channels can be observed accordingly. These planes all belong to the same 3D object.
Figure 3: The TriNeRFLet framework learns features in wavelet domain, which are transformed into feature domain to render 3D objects.
@@ 2 Related Works and Background
$\hat{\bm{C}}=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))\bm{c}_{i},\\ \textrm{where}\quad T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right).\end{aligned}$ (1)
@@
INGP [28] is an important development, which managed to decrease NeRF runtime significantly while getting even better reconstruction performance. It achieved that by using a multiscale hash grid structure alongside a small shallow MLP that represents the scene instead of the big deep one used in previous works. Furthermore, this method introduced new rendering techniques and low-level optimizations in order to get a faster runtime. The multiscale concept of INGP will also be used in our proposed TriNeRFLet.
3D Gaussian Splatting [20] is a state-of-the-art 3D reconstruction approach that takes a different strategy than NeRF. Instead of using an implicit neural function, it represents the scene as a group of 3D Gaussians, and learns their parameters. They introduce a new method for efficiently rendering these Gaussians, thus, managing to train rapidly and achieve 3D rendering in real-time. In run-time, one needs to distinguish between train-time and rendering time. The train time of this method is similar to INGP but its rendering is faster than current NeRF solutions.
Triplane [7, 4] is a NeRF variant, in which the frequency encoding part is replaced by three perpendicular 2D feature planes of dimension $N\times N$ with $C$ channels. To get the feature vector of a point $\bm{x}$, the point is projected onto each of these planes, then the sampled feature values of the projected points are gathered together to form a feature vector. And the final feature vector is obtained by concatenating this feature vector with the frequency encoding of the direction vector $\bm{d}$. These feature planes are learned alongside the MLP weights with a similar approach to what was explained earlier. This process is illustrated in Figure 2(a).
The advantage of Triplane compared to other alternatives is its use of 2D planes, which eases the use of standard 2D networks with it. This has been used in various applications. Bahat et al. [4] utilized it to train a NeRF super-resolution network from multiview training data. By exploiting the triplane structure, a 2D super resolution neural network was trained to perform super resolution on low resolution planes and in this way increase the overall NeRF resolution. Another example is [37], in which a diffusion model was trained to generate a Triplane, thus utilizing the high generalization power of diffusion models and bringing it to 3D generation. Similarly, Triplane is predicted by a transformer-based neural network in [21] in order to have fast text to 3D generation.
@@
Rho et al. [32] implemented a multiscale 3D grid using a grid decomposition method that transforms 2D planes alongside 1D vectors (unrelated to the 2D planes) into a 3D grid, where they used a 2D multiscale wavelet representation to represent the 2D planes. Despite the use of 2D multiscale representation, this method is constrained by 3D grid limitations, namely, it consumes a lot of memory and is limited to low resolutions only. In this work, we deploy the 2D wavelet multiscale representation on Triplane planes directly, thus being freed from 3D grid limitations with a memory-efficient and high-resolution representation.
Diffusion Models [9, 16] have become a very popular tool in recent years for image manipulation and reconstruction tasks [33, 19, 18, 41, 35], where a denoising network is trained to learn the prior distribution of the data. Diffusing models can be used also beyond generation, where given a degraded image, some conditioning mechanism is combined with the learned prior to solve different tasks [3, 2, 8]. For example, in [41, 35, 1] diffusion models were utilized for the problems of deblurring and super-resolution. In this work, we use the Stable Diffusion (SD) upscaler [33] to perform super-resolution on 2D projections of the low-resolution NeRF in order to create a high-resolution version of the NeRF. While the SD upscaler is conditioned both on the low-resolution image and a given text prompt, we use it without text, as explained later in the paper.
3 TriNeRFLet
@@
In rendering, TriNeRFLet is similar to regular Triplane, in the sense that given a point it calculate its features by projecting it to the Triplane and then feed the features to a MLP that outputs the color $c_{i}$ and density $\sigma_{i}$ of the point that are then used by the rendering equation (see Figure 2(a)). Yet, instead of holding the Triplane in the feature domain, TriNeRFLet holds it in the Wavelet domain. Thus, in the optimization, instead of optimizing the values of the triplane directly, we optimize the wavelet coefficients of each pixel (see Figure 2(b)). Note that in the regular Triplane scheme, each plane entry impacts only one location in each plane. In TriNeRFLet, it impacts different resolutions in the wavelent representation of the plane (see Figure 2). Figure 2(c) shows some learned planes for the ship object and their corresponding Wavelet coefficients.
Figure 4: Example of the multiscale property of the wavelet representation. On the right, we the same feautre plane at different resolutions. On the left we see the wavelet representation of the plane at the left, which is of the highest resolution. Note that its wavelet representation include the wavelet representations of the lower-resolution planes (the colors corresponds to the matching between each plane and its wavelet representation).
@@
The learned parameters of TriNeRFLet, alongside the MLP weights, are the wavelet coefficients, and they are learned in an end-to-end fashion. The multiscale structure of the wavelet, allows the method to learn fine details whenever it needs, and to have a lower resolution estimate from lower resolution wavelet layers that are affected by nearby values when the area is not covered by the training pixels. The question remains, what should be done with the coefficients of the higher frequencies that are not updated. From wavelet theory [11], we expect the LH, HL and HH channels to be sparse. Thus, we regularize these channels to be sparse over all resolutions by adding an $L_{1}$ regularization for these coefficients to the training loss. This allow us to increase the resolution of the Triplane without having the problem of having features in it that remain random. In TriNeRFLet, also these features get updated. The overall loss of TriNeRFLet is:
@@
$loss=\sum_{i\in TrainingPixels}\left\lVert\bm{C}_{i}-\hat{\bm{C}_{i}}\right\rVert_{2}^{2}\,+\\ \gamma\sum_{l\in resolutions}\left(\left\lVert\textrm{LH}_{l}\right\rVert_{1}+\left\lVert\textrm{HL}_{l}\right\rVert_{1}+\left\lVert\textrm{HH}_{l}\right\rVert_{1}\right),\end{aligned}$ (2)
@@
Figure 5: TriNeRFLet super-resolution. Low-resolution TriNeRFLet renders LR images using the LR version of the wavelet features, and fits them to the given low-resolution images. High-resolution TriNeRFLet renders HR images using all wavelet levels, and fits them to HR images from HR-set.
To further comprehend the significance of this multiscale wavelet structure against regular Triplane, we compare the behaviour of each method when rendering novel views. In regular Triplane, some features will still have their initial random values after training ends. This introduces artifacts when rendering novel views, as demonstrated in the ship example in Figure 1. Note that these artifacts are absent in TriNeRFLet because all the features in the Triplane are updated either by direct training or using a coarser estimate from lower resolution wavelet coefficients (see Figure 2).
This multiscale wavelet approach enjoys many advantages. First, the $L_{1}$ reguliarziation on the wavelet coefficients prevents the method from over-fitting. In regular NeRF, INGP or Triplane, an over-fitting occurs when training for too many iterations. We have found empirically that this does not happen in our approach.
Another advantage is that the multiscale property of the wavelet representation enables increasing the resolution of the TriNeRFLet during training seamlessly. It is possible to continue to train the higher-resolution TriNeRFLet while using the wavelet coefficients of the lower-resolution TriNeRFLet. The wavelet of the lower-resolution plane is a subset of the wavelet of the higher resolution plane (see Figure 4). We use this flexibility to improve the training by learning the wavelet planes in a coarse to fine manner.
The multiscale wavelet structure is not ideal in terms of training runtime due to the fact that the feature planes need to be reconstructed in each training step. Yet, its rendering runtime is quite fast since the feature planes are reconstructed once only and then the runtime is competitive with other well-known fast NeRF schemes like INGP [28].
[…]
Algorithm 1 TriNeRFLet Super-Resolution
1:Ground truth low resolution (LR) images $\left\{{\bm{x}_{i}}^{gt}_{LR}\right\}_{i=1}^{k}$, LR only steps $s_{low\_res}$, Total steps $s$, Wavelet LR levels $L_{LR}$, Refresh steps $s_{refresh}$, Diffusion step limits $T_{min}\;,\;T_{max}$.

### v2

# TriNeRFLet: A Wavelet Based Triplane NeRF Representation

@@ Abstract
In recent years, the neural radiance field (NeRF) model has gained popularity due to its ability to recover complex 3D scenes. Following its success, many approaches proposed different NeRF representations in order to further improve both runtime and performance. One such example is Triplane, in which NeRF is represented using three 2D feature planes. This enables easily using existing 2D neural networks in this framework, e.g., to generate the three planes. Despite its advantage, the triplane representation lagged behind in 3D recovery quality compared to NeRF solutions. In this work, we propose the TriNeRFLet framework, where we learn the wavelet representation of the triplane and regularize it. This approach has multiple advantages: (i) it allows information sharing across scales and regularization of high frequencies; (ii) it facilitates performing learning in a multi-scale fashion; and (iii) it provides a ‘natural’ framework for performing NeRF super-resolution (SR), such that the low-resolution wavelet coefficients are computed from the provided low-resolution multi-view images and the high frequencies are acquired under the guidance of a pre-trained 2D diffusion model. We show the SR approach’s advantage on both Blender and LLFF datasets.
Keywords:
Neural Radiance Fields (NeRF) Wavelet Multiscale representation 3D Super-Resolution Diffusion Models
@@ 1 Introduction
Figure 1: Our approach improves the quality of NeRF reconstruction (zoom-in). From left to right: TriNeRFLet improvement over Triplane, TriNeRFLet compared to INGP, TriNeRFLet SR improvement, TriNeRFLet SR compared to NeRF-SR.
@@
While being useful due to its special 2D structure, the reconstruction quality achieved by Triplane lagged behind other efficient multiview reconstruction methods such as instant NGP (INGP) [29], which is an improvement of NeRF, and 3D Gaussian splatting [21]. Due to its structure, only Triplane entries that are included in the training views rays are learned. Thus, there might be entries at the planes that will still have their initial random values also after training finishes. These random features are used by novel views that rely on them and therefore they deteriorate the quality of their generation and create artifacts in these novel views.
@@
To tackle these drawbacks, we present a novel Triplane representation that relies on a multiscale 2D wavelet structure. Instead of learning the feature planes directly, we learn the wavelet features of several resolutions, while regularizing the wavelet representation to be sparse. Thus, regions that are covered by the training views are learned in a similar way to the case of a regular Triplane, and regions that are not covered by the training views are updated by a lower resolution estimate according to how much they are affected by their neighboring regions. Figure 1(left) shows the improvement in reconstruction quality attained by this representation. Figure 2(a) illustrates how regions on the plane may affect each other due to the wavelet representation.
@@
In our optimization, we use $L_{1}$ regularization on the wavelet coefficients. The goal is to sparsify the coefficients that are not trained by the given views such that only the coefficients that were trained will impact the reconstruction also from other views. Using $L_{1}$ regularization is inspired by the wavelet literature, which indicates that wavelet features can represent data accurately while being sparse [11].
@@
To further improve the training, we perform it in a multi-scale fashion, starting with a lower resolution version that updates only the coarse coefficients of the wavelet. Then, after some iterations, we increase the resolution and add additional wavelet layers to the representation.
@@
We apply the proposed TriNeRFLet framework to two tasks. First, for 3D reconstruction, where it closes the performance gap of Triplane against other multiview 3D reconstruction methods and makes it competitive with current state-of-the-art (SOTA) methods. Second, we combine it with a pre-trained 2D super-resolution (SR) diffusion models [34, 35] to perform NeRF SR. Figure 1(right) demonstrates the improvement this approach attains. We show in various experiments the advantage of our TriNeRFLet in novel view reconstruction compared to INGP and Triplane and the superiority of our SR solution compared to competing approaches that do not require dedicated training for multiview or 3D data.
@@
(a) The multiscale wavelet representation transfers information between points on the triplane planes: Although the red dot on the plane was not covered by the training images views, it will get a coarser estimate from a lower resolution wavelet layer that corresponds also to the blue dot, which was updated in training.
(b) Example of the multiscale property of the wavelet representation. On the left, we present the wavelet representation. On the right, we show the feature plane of the wavelet representation on the left at different resolutions. Note that the wavelet representation of a higher resolution includes the wavelet representations of the lower-resolution planes (the colors correspond to the matching between each plane and its wavelet representation).
Figure 2: Illustartion of the multiscale property of the wavelet representation.
@@
(b) 3D reconstruction training scheme. First, wavelet features are transformed into the Triplane domain. Next, pixels are rendered using these features in order to fit them to their ground-truth values. The high frequencies channels of LH, HL and HH from all wavelet levels get regularized by $L_{1}$ loss.
(c) In the wavelet domain, the low-frequency coefficients $\textrm{LL}_{N_{base}}$ are observed at the top left corner (the bright spot). The rest are the high frequencies channels $\textrm{LH}_{N_{base}}$, $\textrm{HL}_{N_{base}}$ and $\textrm{HH}_{N_{base}}$. All the other LH, HL and HH channels can be observed accordingly. All these planes belong to the same 3D object.
Figure 4: The TriNeRFLet reconstruction framework learns features in wavelet domain, which are transformed into feature domain to render 3D objects.
@@ 2 Related Works and Background
$\hat{\bm{C}}=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))\bm{c}_{i},\quad\textrm{where}\quad T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right).\end{aligned}$ (1)
@@
INGP [29] is an important development, which managed to decrease NeRF runtime significantly while getting even better reconstruction performance. It achieved that by using a multiscale hash grid structure alongside a small shallow MLP that represents the scene instead of the big deep one used in previous works. Furthermore, this method introduced new rendering techniques and low-level optimizations in order to get a faster runtime. The multiscale concept of INGP will also be used in our proposed TriNeRFLet framework.
3D Gaussian Splatting [21] is a state-of-the-art 3D reconstruction approach that takes a different strategy than NeRF. Instead of using an implicit neural function, it represents the scene as a group of 3D Gaussians, and learns their parameters. It introduces a new method for efficiently rendering these Gaussians, thus, managing to train rapidly and achieving 3D rendering in real-time. Note that one needs to distinguish between training and rendering time. While its training time is similar to INGP, its rendering is faster than NeRF solutions.
Triplane [7, 4] is a NeRF variant, in which the frequency encoding part is replaced by three perpendicular 2D feature planes of dimension $N\times N$ with $C$ channels. To get the feature vector of a point $\bm{x}$, the point is projected onto each of these planes, then the sampled feature values of the projected points are gathered together to form a feature vector. The final feature vector is obtained by concatenating this feature vector with the frequency encoding of the direction vector $\bm{d}$. These feature planes are learned alongside the MLP weights with a similar approach to what was explained earlier. Fig. 4(a) illustrates this process.
The advantage of Triplane compared to other alternatives is its use of 2D planes, which eases the use of standard 2D networks with it. This has been used in various applications. Bahat et al. [4] utilized it to train a NeRF super-resolution (SR) network from multiview training data. By exploiting the triplane structure, a 2D SR neural network was trained to perform SR on low resolution planes and in this way increase the overall NeRF resolution. Another example is [38], in which a diffusion model was trained to generate a Triplane. It utilizes the generalization power of diffusion models and brings them to 3D generation. Similarly, Triplane is predicted by a transformer-based neural network in [22] in order to have fast text to 3D generation.
@@
Rho et al. [33] implemented a multiscale 3D grid using a grid decomposition method that transforms 2D planes alongside 1D vectors (unrelated to the 2D planes) into a 3D grid, where they used a 2D multiscale wavelet representation to represent the 2D planes. Despite the use of 2D multiscale representation, this method is different in essence from TriNeRFLet, which deploys the 2D wavelet multiscale representation on Triplane planes directly, thus being freed from 3D grid limitations.
Diffusion Models [9, 16] have become a very popular tool in recent years for image manipulation and reconstruction tasks [34, 20, 19, 42, 36], where a denoising network is trained to learn the prior distribution of the data. Diffusing models can be used also beyond generation, where given a degraded image, some conditioning mechanism is combined with the learned prior to solve different tasks [3, 2, 8]. For example, in [42, 36, 1] diffusion models were utilized for the problems of deblurring and SR. In this work, we use the Stable Diffusion (SD) upscaler [34] to perform SR on 2D projections of the low-resolution NeRF in order to create a high-resolution version of the NeRF. While the SD upscaler is conditioned both on the low-resolution image and a given text prompt, we use it without text, as explained later in the paper.
The authors of [43] proposed to fine-tune a diffusion model on the views of a given scene to regularize unseen views. In a similar spirit, we use images generated by a diffusion model to fit the scene and thus improve NeRF SR.
3 TriNeRFLet Framework
@@
In rendering, TriNeRFLet is similar to regular Triplane, in the sense that given a point it calculate its features by projecting it to the Triplane and then feed the features to a MLP that outputs the color $c_{i}$ and density $\sigma_{i}$ of the point that are then used by the rendering equation (see Figure 4(a)). Yet, instead of holding the Triplane in the feature domain, TriNeRFLet holds it in the Wavelet domain and optimizes the wavelet coefficients. Note that in the regular Triplane scheme, each plane entry impacts only one location in each plane. In TriNeRFLet, it impacts different resolutions in the wavelent representation of the plane (see Figure 2(a)). Figure 4(c) shows some learned planes for the ship object and their corresponding Wavelet coefficients.
@@
The learned parameters of TriNeRFLet, alongside the MLP weights, are the wavelet coefficients, and they are learned in an end-to-end fashion. The multiscale structure of the wavelet, allows the method to learn fine details whenever it needs, and to have a lower resolution estimate from lower resolution wavelet layers that are affected by nearby values when the area is not covered by the training pixels. The question remains, what should be done with the coefficients of the higher frequencies that are not updated.
Wavelet regularization. From wavelet theory [11], we expect the LH, HL and HH channels to be sparse. Thus, we regularize these channels to be sparse over all resolutions by adding an $L_{1}$ regularization for these coefficients to the training loss. This allows us to increase the resolution of the planes without having the problem of having features in it that remain random or redundant. In TriNeRFLet, also these features get updated.
To further comprehend the significance of the multiscale wavelet structure and the regularization we use for it, we compare the behavior to rendering novel views with a regular Triplane. In a regular Triplane, some features will still have their initial random values after training ends. This introduces artifacts when rendering novel views, as demonstrated in the ship example in Figure 1. Note that these artifacts are absent in TriNeRFLet because all the features in the planes are updated either by direct training or using a coarser estimate from lower resolution wavelet coefficients (see Figure 2(a)).
Multiscale training. A key advantage of TriNeRFLet is that the multiscale property of the wavelet representation enables increasing the resolution of the TriNeRFLet during training seamlessly. It is possible to continue to train the higher-resolution TriNeRFLet while using the wavelet coefficients of the lower-resolution TriNeRFLet. The wavelet of the lower-resolution plane is a subset of the wavelet of the higher resolution plane (see Figure 2(b)). We use this flexibility to improve the training by learning the wavelet planes in a coarse to fine manner. As we explain later, we leverage the multiscale structure of the wavelet both for NeRF reconstruction and SR.
The multiscale wavelet structure is not ideal in terms of training runtime due to the fact that the feature planes need to be reconstructed in each training step. Yet, its rendering runtime is quite fast since the feature planes are reconstructed just once and then the runtime is competitive with other well-known fast NeRF schemes like INGP [29].
4 TriNeRFLet Applications
4.1 TriNeRFLet Reconstruction
The TriNeRFLet framework is straightforwardly applied to 3D scene reconstruction task. Given the training pixels, wavelet features are optimized end-to-end on the pixel reconstruction loss in addition to the wavelet high frequencies regularization (see Figure 4(b)). The overall loss is:
@@ 4.1 TriNeRFLet Reconstruction
$\mathcal{L}=\sum_{i\in TrainingPixels}\left\lVert\bm{C}_{i}-\hat{\bm{C}_{i}}\right\rVert_{2}^{2}\,+\gamma\sum_{l\in resolutions}\left(\left\lVert\textrm{LH}_{l}\right\rVert_{1}+\left\lVert\textrm{HL}_{l}\right\rVert_{1}+\left\lVert\textrm{HH}_{l}\right\rVert_{1}\right),$
@@
As mentioned before, the flexible multiscale structure of TriNeRFLet enables the training to be in a coarse-to-fine manner. This mitigates the runtime overhead of the wavelet inverse transform. At the beginning of the training, the method usually learns high-level features, and then when it starts to converge it learns the low-level details, which is coherent with the coarse-to-fine approach.
4.2 TriNeRFLet Super-Resolution (SR)
We turn to present a novel NeRF SR approach that relies on the proposed multiscale structure. This method employs a 2D pre-trained stable diffusion (SD) upscaler [34] for SR “guidance”. Our approach does not require any kind of low-high resolution 3D scene pairs for supervision as done in other works [4]. Utilizing the robustness of the SD upscaler, our approach has the potential to handle different types of scenes and a range of resolutions.
Given a set of low-resolution (LR) multiview images of a given scene, the goal is to generate high-resolution (HR) views of this scene. We first apply TriNeRFLet using the LR images, providing a LR 3D reconstruction. In this reconstruction, we learn only the low frequencies of the wavelet representation. Denote the size of the LR planes by $N_{LR}$ and the number of wavelet levels by $L_{LR}$. For the HR planes, denote the size by $N$ and the number of levels by $L$. Due to the multiscale nature wavelet representations, in the HR TriNeRFLet we just need to learn the high frequencies of the wavelet, i.e., the first $L_{LR}$ levels of the HR wavelet planes are shared with the LR TriNeRFLet reconstruction.
Algorithm 1 TriNeRFLet Super-Resolution (SR)
1:Ground truth low resolution (LR) images $\left\{{\bm{x}_{i}}^{gt}_{LR}\right\}_{i=1}^{k}$, LR only steps $s_{LR}$, Total steps $s$, Wavelet LR levels $L_{LR}$, Refresh steps $s_{refresh}$, Diffusion step limits $T_{min}\;,\;T_{max}$.